OpenStudy (anonymous):

Which is the equation of a line that passes through the point (3, -4) and has a slope of 2?

OpenStudy (rajee_sam):

y-(-4) = 2 (x-3) y+4 = 2x - 6 y = 2x -6 - 4 y = 2x - 10 : Ans.